OpenStudy (anonymous):

Solve for x: 3x − 5 = 2x + 6 1 −1 11 −11

OpenStudy (ivettef365):

you need to have the x on one side and integers on the other, switch signs when switching numbers as follows: 3x -2x = 6 + 5

OpenStudy (dan815):

http://www.youtube.com/watch?feature=endscreen&v=KsgWLV3RiRQ&NR=1 some education

OpenStudy (dan815):

edumacation

OpenStudy (mathstudent55):

First, subtract 2x from both sides to get all x-terms on the left side. Then add 5 to both sides to get all number terms on the right side. Then divide both sides by the number multiplying x to find the value of x.

OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution